# LinkPrompt: Natural and Universal Adversarial Attacks on Prompt-based Language Models

### v1

@@ 1 Introduction
Prompt-based learning is a new language model training paradigm that aims to adapt the Pre-trained Language Models (PLMs) to perform well on the downstream tasks, which revitalizes the performance benchmarks across various natural language processing (NLP) tasks Radford et al. (2019); Brown et al. (2020); Schick and Schütze (2020). By equipping input sentences with designed prompt templates Liu et al. (2023), prompt-based learning converts a text classification task into a next-word prediction task. Then the PLMs are fine-tuned under the prompt-based learning framework to get Prompt-based Fine-tuned Models (PFMs) that are specific to downstream tasks. Specialized prompt can effectively connect PLMs with downstream tasks in few-shot scenarios Winata et al. (2021); Tsimpoukelli et al. (2021). The process of prompt-based learning is demonstrated in Figure 1.
@@
To further enhance the performance of PLMs and PFMs, instead of using a fixed prompt template to fine-tune the model, some methods are proposed to optimize the prompts by maximizing the prediction outcomes. For example, AutoPrompt Shin et al. (2020) applied a gradient-based search strategy to optimize a universal prompt template with a fixed length of tokens specific to a downstream task, thus improving the model training efficiency and the generalization ability.
@@
Similar to the adversarial attack on simple text classification models, PLMs under prompt-based learning frameworks also suffer from potential adversarial threats. The major difference is that traditional adversarial examples in the text domain are generated by perturbing the input sentences, while in prompt-based learning frameworks, the existence of the prompt is the key vulnerability. Wallace et al. (2019) first propose a universal adversarial attack on PLMs by optimizing universal adversarial triggers (UATs) that can cause a model to give wrong predictions to any inputs. Xu et al. (2022) proposed a universal adversarial attack named AToP under the prompt-based learning paradigm and proved that the adversarial trigger optimized to target the PLMs can also transfer to the PFMs.
Although AToP can successfully diminish the prediction accuracy of PFMs, such UATs have a limitation in naturalness, which means they are meaningless combinations of tokens and symbols that can be easily detected by adaptive defense techniques with simple heuristics. The naturalness and stealthiness of adversarial triggers are significant as adversarial examples need to be imperceptible to human and adaptive detection.
To generate more powerful and natural adversarial triggers, we introduce a universal adversarial attack algorithm named LinkPrompt, which can not only fool the prompt-based learned language model into making wrong predictions but also maintain the naturalness among the generated adversarial triggers. Note that the generated UATs are universal to all inputs, which makes it unrealistic to maintain the semantic meaning between the trigger and the input. Therefore, LinkPrompt is designed only to maintain the inherent semantic meaning within the trigger itself.
The process of LinkPrompt attack can be described in two phases. The first phase is trigger selection, where we optimize the trigger tokens through a large text corpus (e.g. Wikitext, Merity et al., 2016) on PLMs. Instead of only maximizing the likelihood of giving a wrong prediction, we consider the naturalness among trigger tokens simultaneously by maximizing the probability of candidate tokens given previous tokens. Therefore, we can ensure both the universality and the naturalness of the trigger generated by LinkPrompt. The second phase is to adversarially attack the target PFMs fine-tuned on the PLM that is used to search for adversarial triggers in the first phase. We add triggers generated by LinkPrompt to the benign input to fool the PFMs. The illustration of these two phases is demonstrated in Figure 2.
@@
  • We leverage the angle-optimized text embedding model AnglE Li and Li (2023) and ChatGPT OpenAI (2024) as additional evaluation methods than perplexity to better measure the naturalness of UATs generated by LinkPrompt.
  • We conduct the transferability study of LinkPrompt on BERT Devlin et al. (2018), as well as on an open-sourced large language model Llama2 Touvron et al. (2023) and an API-accessed large language model GPT-3.5-turbo.
@@ 3.3 Trigger Selection
Semantic objectives. To achieve the semantic goal which is to maintain the semantic meaning among the adversarial trigger tokens, the second objective is to maximize the probability of the current candidate token given the previous tokens. Leveraging the predictive ability of the PLMs, such prediction probability can reflect the semantic relevance between the candidate token and the preceding context. To maximize the inherent semantic naturalness of a specific trigger $\mathbf{t}$ of length $L$, we use the probability of the current candidate token $t_{i}$ being predicted based on the previous tokens to represent the semantic naturalness between the current token with the previous tokens. Note that we want to maximize the prediction probability which equals minimizing the semantic loss. Therefore, the loss can be defined as the negative of the average prediction probability of each token given the previous token in the trigger:
@@ 4 Experiment
In this section, we first introduce the configuration of our experiments, including the victim model, datasets, prompt templates, baseline, and evaluation metrics. Then, we evaluate the effectiveness and naturalness of UATs generated by LinkPrompt. Following that, we demonstrate the transferability of LinkPrompt on Bert, Llama2, and GPT-3.5-turbo. Finally, we propose an adaptive defense and demonstrate the stability of LinkPrompt.
@@ 4.1 Configurations
Baseline and evaluation metrics. We compare LinkPrompt with AToP, a state-of-the-art universal adversarial attack on PFM. The objective of AToP is the first loss term of Equation 3, which is equivalent to the situation that $\alpha$ is equal to 0.
@@
Last, the Semantic Similarity Score (SSS) represents the semantic similarity between the original and modified sentences. The assumption is that the more similar the adversarially perturbed sentence is to the original sentence, the more naturalness the UAT maintains, and the less it is suspicious to the adversarial detection. To measure SSS, We use AnglE, an angle-optimized text embedding model specialized for semantic textual similarity tasks, to obtain and compare the embedding distance as shown in Figure 3. The similarity score can be calculated as $\text{sim}(\mathbf{u},\mathbf{v})=1-\arccos\left(\frac{\mathbf{u}\cdot\mathbf{v}}{\|\mathbf{u}\|\|\mathbf{v}\|}\right)/\pi$ Cer et al. (2018), where $\mathbf{u}$ and $\mathbf{v}$ present the embedding of perturbed sentence and original sentence respectively. Higher SSS indicates higher semantic similarity.
@@ 4.3 UATs Naturalness Evaluation
Semantic Similarity Score. The effectiveness and naturalness of generated UATs are controlled by the weight $\alpha$ to balance the two loss terms. It is obvious that a greater $\alpha$ will push the optimization process to generate more natural UATs while suffering the trade-off on the ASR, and vice versa. Therefore, we plot the trade-off between the attack effectiveness and UAT naturalness with ASR and semantic similarity score (SSS) in Figure 5. We can note that UATs generated by LinkPrompt are gathered on the right-upper part of each plot, which indicates that LinkPrompt can achieve comparable ASRs while having higher SSS.
Evaluation by ChatGPT. To analyze the naturalness of UATs generated by LinkPrompt from the perspective of advanced language models, we involve ChatGPT in our measurement. It is worth noting that, unlike using sentence encoders for naturalness assessment, we need to avoid triggers affecting the model’s understanding of sentence meaning. Therefore, we form two statements for one input by appending it with our UATs and those in AToP respectively, and then use ChatGPT to determine which statement is more natural. The prompt used for measurement can be found in Appendix C.1. We use the winning rate of 300 sentences in each dataset to quantify the evaluation and the comparison results are shown in Table 2. The high winning rate of LinkPrompt further demonstrates its naturalness.
@@ 5 Related Articles Sega Console Controller Yourselves Share Skip Disable JavaScript
Triggers Visualization. We further visualize the UATs generated by LinkPrompt to demonstrate the naturalness. Table 3 captures the triggers found by both LinkPrompt and the baseline AToP under different trigger lengths. There are almost no meaningless symbols in LinkPrompt and the higher semantic relevance between the tokens can be observed.
@@ 4.5 Transferability
The UATs we evaluated in the previous sections are generated on RoBERTa-large. The transferability is crucial to adversarial perturbations which indicates the generalization ability of the generated UATs. Therefore, in this section, we want to evaluate whether the triggers we found on RoBERTa-large can lead to misclassification to other PFMs regardless of their structure.
@@
Len Matrices $\boldsymbol{\alpha}$ AG SST2 IMDB HATE FN FR
@@
Transfer to GPT-3.5. In addition to attacking open-source large language model, we also conduct preliminary verification on API-based black-box large language model GPT-3.5-turbo. Due to the inability to perform prompt-based fine-tuning on black-box models, we directly incorporate UATs into input sentences and use the same prompts in Llama2 experiment for testing. The test results are shown in Table 5. It can be observed that UATs also achieve good attack effectiveness on black-box language model. We speculate that this is because UATs influence the model’s understanding of sentence meaning, and this influence is widespread across various architectures of language models.
@@ 4.6 Adaptive Defense
We further propose a perplexity filter as an adaptive defense against LinkPrompt. Although LinkPrompt can maintain the semantic naturalness within the UAT, it is still irrelevant to the input sentences for the universality. Therefore, we proposed a perplexity detection filter inspired by ONION Qi et al. (2020) to test the stealthiness of UATs generated by LinkPrompt.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 5 Conclusion
We propose LinkPrompt, a universal adversarial attack algorithm on PFMs that can not only mislead the PFMs to give wrong predictions but also maintain naturalness. Compared with previous work, LinkPrompt can achieve a higher attack success rate while increasing the naturalness of triggers as well. We also evaluate the transferability of LinkPrompt to different model structures. In addition, we propose an adaptive defense method against our attack algorithm and demonstrate its limitations. In further study, we will explore new methods to generate triggers that are more stealthy with the assistance of large language models. It is also worthwhile to transfer such a method to other tasks or larger models.
@@
In this paper, we act as an attacker and propose an algorithm to generate UATs that are both effective and natural, which also have strong transferability and stability. It is possible that the UATs or our method are being maliciously used in terms of attacking existing language models. However, we consider research on such attacks to be significant to improve the robustness of state-of-the-art large language models and we intend to release both the algorithm and the generated triggers so that better defense can be developed in the future. In addition, we can gain insights from our experimental findings, resulting in a better understanding of the prompt-based fine-tuning paradigm and the language models as well.
@@
We conclude the limitations of our work in three aspects: First, to maintain the universality and effectiveness of triggers, which means that they can be adapted to any PFMs and inputs while having a high ASR, the triggers generated by LinkPrompt are still not natural enough in human evaluation. This may be explained by that there exists an inherent trade-off either between the universality or the performance and the fluency of triggers, which has also been proved in previous works. To improve the triggers’ naturalness in the human evaluation system, developing the adversarial attack algorithm combined with techniques such as Reinforcement Learning from Human Feedback (RLHF) can be a potential solution. Second, in this paper, we mainly focus on the classification tasks and choose the masked language model RoBERTa-large as our victim model due to its good performance in such tasks. However, PFMs specific to generation tasks such as translation and dialogue can also suffer from adversarial threats. It is worthwhile to expand LinkPrompt to other tasks and larger-scale language models. Third, the adaptive defense based on a unified perplexity filter does not work well on LinkPrompt, which can be evidenced by the increase in ASR of certain triggers and the significant decrease in accuracy. In further studies, we intend to propose a stronger defense against LinkPrompt with the assistance of large language models. Instead of just computing the perplexity of a sentence, we can train a language model to determine whether a sentence is semantically natural or not.
@@ B.1 Attack results of LinkPrompt on RoBERTa-large
Refer to caption ((a)) ASR results of 3-token triggers. Refer to caption ((b)) ASR results of 7-token triggers.
@@ B.2 Attack results of LinkPrompt on Llama2
Refer to caption ((a)) ASR results of 3-token triggers. Refer to caption ((b)) ASR results of 7-token triggers.

### v2

@@ 1 Introduction
Prompt-based learning is a new language model training paradigm that aims to adapt the Pre-trained Language Models (PLMs) to perform well on the downstream tasks, which revitalizes the performance benchmarks across various natural language processing (NLP) tasks Radford et al. (2019); Brown et al. (2020); Schick and Schütze (2020). By equipping input sentences with designed prompt templates Liu et al. (2023), prompt-based learning converts a text classification task into a next-word prediction task. Then the PLMs are fine-tuned under the prompt-based learning framework to get Prompt-based Fine-tuned Models (PFMs) that are specific to downstream tasks. Specialized prompts can effectively connect PLMs with downstream tasks in few-shot scenarios Winata et al. (2021); Tsimpoukelli et al. (2021). The process of prompt-based learning is demonstrated in Figure 1.
@@
To further enhance the performance of PLMs and PFMs, instead of using a fixed prompt template to fine-tune the model, some methods are proposed to optimize the prompts by maximizing the probability of desired outcomes. For example, AutoPrompt Shin et al. (2020) applied a gradient-based search strategy to optimize a universal prompt template with a fixed length of tokens specific to a downstream task, thus improving the model training efficiency and the generalization ability.
@@
Similar to the adversarial attack on simple text classification models, PLMs under prompt-based learning frameworks also suffer from potential adversarial threats. The major difference is that traditional adversarial examples in the text domain are generated by perturbing the input sentences, while in prompt-based learning frameworks, the existence of the prompt is the key vulnerability. Wallace et al. (2019) first propose a universal adversarial attack on PLMs by optimizing universal adversarial triggers (UATs) that can cause a model to give wrong predictions to any inputs.
Furthermore, recent studies discovered that the vulnerability to adversarial attacks of PLMs can also be carried to the PFMs under the prompt-based learning paradigm. Xu et al. (2022) proposed a universal adversarial attack named AToP under the prompt-based learning paradigm and proved that the adversarial trigger optimized to target the PLMs can also transfer to the PFMs. Although AToP can successfully diminish the prediction accuracy of PFMs, such UATs have a limitation in naturalness, which means they are meaningless combinations of tokens and symbols that adaptive defense techniques with simple heuristics can easily detect.
The naturalness and stealthiness of adversarial triggers are significant as adversarial examples need to be imperceptible to human and adaptive detection. To generate more powerful and natural adversarial triggers, we introduce a universal adversarial attack algorithm named LinkPrompt, which can not only fool the prompt-based fine-tuned language model into making wrong predictions but also maintain the naturalness among the generated adversarial triggers. Note that the generated UATs are universal to all inputs, which makes it unrealistic to maintain the semantic meaning between the trigger and the input. Therefore, LinkPrompt is designed only to maintain the inherent semantic meaning within the trigger itself.
The process of LinkPrompt attack can be described in two phases. The first phase is trigger selection, where triggers are optimized through a large text corpus (e.g. Wikitext, Merity et al., 2016) on a PLM. Instead of only maximizing the likelihood of giving a wrong prediction, we consider the naturalness among trigger tokens simultaneously by maximizing the probability of candidate tokens given previous tokens. Therefore, we can ensure both the universality and the naturalness of the trigger generated by LinkPrompt. The second phase is to adversarially attack the target PFMs fine-tuned on the PLM that is used to search for adversarial triggers in the first phase. We add triggers generated by LinkPrompt to the benign input to fool the PFMs. The illustration of these two phases is demonstrated in Figure 2.
@@
  • We leverage an Angle-optimized text Embedding model called AnglE Li and Li (2023) and ChatGPT OpenAI (2024) as additional evaluation methods than perplexity to better measure the naturalness of UATs generated by LinkPrompt.
  • We conduct the transferability study of LinkPrompt on BERT Devlin et al. (2018), as well as on an open-sourced large language model Llama2 Touvron et al. (2023) and an API-accessed large language model GPT-3.5-turbo OpenAI (2024).
@@ 3.3 Trigger Selection
Semantic objectives. To achieve the semantic goal which is to maintain the semantic meaning among the adversarial trigger tokens, the second objective $\mathcal{L}_{sem}$ is to maximize the probability of the current candidate token given the previous tokens. Leveraging the predictive ability of the PLMs, such prediction probability can reflect the semantic relevance between the candidate token and the preceding context. To maximize the inherent semantic naturalness of a specific trigger $\mathbf{t}$ of length $L$, we use the probability of the current candidate token $t_{i}$ being predicted based on the previous tokens to represent the semantic naturalness between the current token with the previous tokens. In other words, we want to maximize the average prediction probability of each token given the previous token in the trigger, which equals to minimize the following loss:
@@ 4 Experiment
In this section, we first introduce the configurations of our experiments, including the victim model, datasets, prompt templates, baseline, and evaluation metrics. Then, we evaluate the effectiveness and naturalness of UATs generated by LinkPrompt. Following that, we demonstrate the transferability of LinkPrompt on Bert, Llama2, and GPT-3.5-turbo. Finally, we propose an adaptive defense and demonstrate the stability of LinkPrompt.
@@ 4.1 Configurations
Baseline and evaluation metrics. We compare LinkPrompt with AToP, a state-of-the-art universal adversarial attack on PFM. The objective of AToP is the first loss term of Equation 3, which is equivalent to the situation that $\alpha$ equals 0.
@@
Last, the Semantic Similarity Score (SSS) represents the semantic similarity between the original and modified sentences. The assumption is that the more similar the adversarially perturbed sentence is to the original sentence, the more naturalness the UAT maintains, and the less it is suspicious to the adversarial detection. To measure SSS, We use AnglE Li and Li (2023), an angle-optimized text embedding model that achieves state-of-the-art performances in semantic textual similarity tasks, to obtain the embedding distance as shown in Figure 3. The similarity score can be calculated as $\text{sim}(\mathbf{u},\mathbf{v})=1-\arccos\left(\frac{\mathbf{u}\cdot\mathbf{v}}{\|\mathbf{u}\|\|\mathbf{v}\|}\right)/\pi$ Cer et al. (2018), where $\mathbf{u}$ and $\mathbf{v}$ present the embedding of perturbed sentence and original sentence respectively. Higher SSS indicates higher semantic similarity.
@@ 4.3 UATs Naturalness Evaluation
Semantic Similarity Score. The effectiveness and naturalness of generated UATs are controlled by the weight $\alpha$ to balance two loss terms. A greater $\alpha$ will push the optimization to generate more natural UATs while suffering the trade-off on the ASR, and vice versa. Therefore, we plot the trade-off between the attack effectiveness and UAT naturalness with ASR and semantic similarity score (SSS) in Figure 5. We can note that UATs generated by LinkPrompt are gathered on the right-upper part of each plot, which indicates that LinkPrompt can achieve comparable ASRs while having higher SSS.
Evaluation by ChatGPT. To further demonstrate the naturalness of UATs generated by LinkPrompt, we incorporate ChatGPT as another measurement. We generate two statements for each input by appending it with UATs optimized by LinkPrompt and AToP respectively, and then use ChatGPT to determine which statement is more natural. The prompt used for measurement can be found in Appendix C.1. We demonstrate the winning rate of LinkPrompt on 300 input queries in each dataset in Table 2. If the winning rate exceeds 50%, LinkPrompt exhibits superior naturalness compared to AToP. As shown in Table 2, the winning rate of all datasets surpasses 50% and in some cases of SST2 and HATE datasets even achieves 90%, which proves LinkPrompt significantly enhances UATs’ naturalness.
@@ 5 Related Articles Sega Console Controller Yourselves Share Skip Disable JavaScript
Triggers Visualization. We further visualize the UATs generated by LinkPrompt to demonstrate the naturalness. Table 3 captures the triggers found by both LinkPrompt and the baseline AToP under different trigger lengths. There are hardly meaningless symbols in LinkPrompt and the higher semantic relevance between the tokens can be observed.
@@ 4.5 Transferability
The UATs we evaluated in the previous sections are generated on RoBERTa-large. The transferability is crucial to adversarial perturbations which indicates the generalization ability of UATs. Therefore, in this section, we evaluate whether the triggers optimized on RoBERTa-large can lead to misclassification to other PFMs regardless of their structure.
@@
Len Metrics $\boldsymbol{\alpha}$ AG SST2 IMDB HATE FN FR
@@
Transfer to GPT-3.5. In addition to testing the transferability of LinkPrompt on open-source large language models, we also conduct preliminary verification on API-based black-box large language model GPT-3.5-turbo. Due to the inability to perform prompt-based fine-tuning on black-box models, we directly incorporate UATs into input sentences and use the same setting in the Llama2 experiment for testing. As shown in Table 5, UATs also achieve good attack effectiveness on the black-box language model. We speculate that this is because UATs influence the model’s understanding of sentence meaning, and this influence is widespread across various architectures of language models.
@@ 4.6 Adaptive Defense
We further propose a perplexity filtering as an adaptive defense against LinkPrompt. Although LinkPrompt can maintain the semantic naturalness within the UAT, it is still irrelevant to the input sentences for the universality. Therefore, we propose a perplexity detection filter inspired by ONION Qi et al. (2020) to test the stealthiness of UATs generated by LinkPrompt.
@@
We compare the stealthiness of LinkPrompt with the baseline method AToP on two datasets. We select UATs generated by LinkPrompt that have comparable ASR with AToP to conduct a fair comparison. Table 6 shows the change of ASR after applying the filtering. First, we compare the drop of ASR under different trigger lengths (AToP and $\textit{LinkPrompt}_{avg}$). As shown in Table 6, the drop of ASR ($\Delta$ columns) on LinkPrompt after the filtering is overall lower than AToP on both datasets, except the result on SST-2 with trigger length 7, which indicates that LinkPrompt is more resilient to the perplexity based adaptive defense.
Second, we compare the drop of ASR under different original ASRs (indicated as $\textit{LinkPrompt}_{low}$ and $\textit{LinkPrompt}_{high}$ in Table 6), as the original low and high ASRs have a different trend. Remember we design an objective function with a weighted sum of two loss terms from the attack and the naturalness perspective respectively. We can adjust the weight $\alpha$ to control the naturalness of generated UATs. Generally, a higher $\alpha$ can result in more natural but less successful UATs, and vice versa. In Table 6, ASRs of less effective triggers ($\textit{LinkPrompt}_{low}$) even rise after the process of such a perplexity filter and the accuracy drops heavily on both tasks. This indicates the limitation of such an outlier detecting method towards LinkPrompt.
@@
$\textit{LinkPrompt}_{avg}$-3 28.72 -12.45 39.58 -19.31
$\textit{LinkPrompt}_{high}$-3 31.65 -20.73 35.15 -42.89
$\textit{LinkPrompt}_{low}$-3 25.80 -4.17 44.01 +4.27
@@
$\textit{LinkPrompt}_{avg}$-5 54.77 -20.59 48.30 -7.10
$\textit{LinkPrompt}_{high}$-5 46.07 -53.68 45.14 -27.85
$\textit{LinkPrompt}_{low}$-5 63.47 +12.51 51.46 +13.65
@@
$\textit{LinkPrompt}_{avg}$-7 63.13 +0.52 50.04 -9.71
$\textit{LinkPrompt}_{high}$-7 58.07 -15.94 50.83 -25.34
$\textit{LinkPrompt}_{low}$-7 68.18 +16.98 49.25 +5.93
@@ 5 Conclusion
We propose LinkPrompt, a universal adversarial attack algorithm on PFMs that can not only mislead the PFMs to give wrong predictions but also maintain naturalness. Compared with previous work, LinkPrompt can achieve a higher attack success rate while increasing the naturalness of triggers as well. We also evaluate the transferability of LinkPrompt to different model structures. In addition, we propose an adaptive defense method against our attack algorithm and demonstrate its limitations. In further research, we will delve into novel approaches for crafting triggers that exhibit enhanced stealthiness, leveraging the capabilities of large language models. Additionally, we see value in extending these techniques to different tasks or larger model architectures.
Acknowledgment
We thank all reviewers for their constructive comments. This work is supported by Shanghai Engineering Research Center of Intelligent Vision and Imaging.
@@ Acknowledgment
In this paper, we act as an attacker and present an algorithm for generating effective and natural Universal Adversarial Triggers (UATs), demonstrating strong transferability and stability as attack vectors. It is possible that the UATs or our method are being maliciously used in terms of attacking existing language models. However, researching such attacks is crucial to bolster the robustness of modern large language models. We aim to release both the algorithm and the generated triggers to aid in the development of improved defense strategies. Additionally, our experimental results contribute valuable insights into the prompt-based fine-tuning paradigm and advance our understanding of language models.
@@
We outline the limitations of our study as follows:
1. The UATs produced by LinkPrompt exhibit deficiencies in naturalness according to human evaluations, highlighting a compromise between universality, performance, and fluency. Enhancing their naturalness may be achievable through the development of adversarial attack algorithms coupled with techniques like Reinforcement Learning from Human Feedback (RLHF).
2. Our investigation predominantly focuses on classification tasks, utilizing the masked language model RoBERTa-large as our target due to its commendable performance in such contexts. Nevertheless, PFMs tailored to generation tasks such as translation and dialogue may also face adversarial vulnerabilities. Expanding LinkPrompt to these tasks and larger-scale language models is worth considering.
3. Our adaptive defense strategy, relying on a unified perplexity filter, demonstrates ineffectiveness against LinkPrompt, as indicated by increased ASR for certain triggers and a notable decline in accuracy. Future endeavors will aim to devise more robust defenses, potentially leveraging large language models to assess sentence semantic naturalness rather than relying solely on perplexity.
@@ B.1 Attack results of LinkPrompt on RoBERTa-large
((a)) ASR results of 3-token triggers.
((b)) ASR results of 7-token triggers.
@@ B.2 Attack results of LinkPrompt on Llama2
((a)) ASR results of 3-token triggers.
((b)) ASR results of 7-token triggers.